OpenStudy (anonymous):

Help! Find the vertices and foci of the ellipse. 25x2 + 81y2 = 81

OpenStudy (anonymous):

\[25x^2 + 81y^2 = 81\]

OpenStudy (amistre64):

might wanna start out by making it =1

OpenStudy (anonymous):

so it would be \[\frac{ 25 }{ 81 } x^2 +y^2 = 1 ?\]

OpenStudy (anonymous):

ya now u compare it with d std. form

OpenStudy (anonymous):

would a^2 = 81? and b^2 = 1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

so a = 9 b=1 c= sqrt(80) how do i find the vertices? (x,9) and (x,-9) ? what is x? same with the x in the foci? would it be 25?

OpenStudy (anonymous):

u can now find foci and do vertices mean the extreme point on major and minor axis???